OpenStudy (anonymous):

14. Which equation is an identity? (1 point)11 – (2v + 3) = –2v – 8 5w + 8 – w = 6w – 2(w – 4) 7m – 2 = 8m + 4 – m 8y + 9 = 8y – 3

terenzreignz (terenzreignz):

You could try simplifying each equation. If it leads to something that's obviously true (say 4 = 4) instead of a concrete solution (like y = 3) or a blatant untruth (like 9 = -3) then it's the identity.

OpenStudy (anonymous):

im not really good with stuff like this

OpenStudy (anonymous):

ive gone to 3 different schools this year and they all teach different things. its annoying

terenzreignz (terenzreignz):

Well, we got to live with what we got :) Tell you what, pick one of your choices, I'll demonstrate simplifying it for you.

OpenStudy (anonymous):

7m – 2 = 8m + 4 – m

terenzreignz (terenzreignz):

Combine like terms on each side. On the left, there are no like terms 7m - 2 (like terms are terms with the exact same set of variables on them) On the right 8m + 4 - m You can combine 8m and -m to get 7m 7m + 4. so you end up with 7m - 2 = 7m + 4 You can add -7m on both sides as per the Addition Property of Equality: 7m - 2 - 7m = 7m + 4 - 7m Combine like terms again: -2 = +4 And you get something which is false. So this is NOT the identity.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

id have to say its this one 11 – (2v + 3) = –2v – 8

OpenStudy (anonymous):

do think it is right

terenzreignz (terenzreignz):

Simplify it, we'll see.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh sorry im back

terenzreignz (terenzreignz):

Sure. Just keep simplifying :)